OpenStudy (anonymous):

@myininaya @Frostbite @sirm3d @samtasticc @whpalmer4 @qaisar900 @calmat01 @ChemicalX @calmat01

myininaya (myininaya):

Yes @danya1 ?

OpenStudy (anonymous):

CAN U HELP ME SIMPLIFY THIS

myininaya (myininaya):

Do you know law of exponents?

OpenStudy (anonymous):

no

myininaya (myininaya):

\[x^n x^m=x^{n+m}\] \[\frac{x^n}{x^m}=x^{n-m}\] \[\frac{1}{x^{-n}}=x^n\] \[x^{-n}=\frac{1}{x^n}\] \[(x^ry^t)^s=x^{rs}y^{ts}\] Here are a few laws... You tell me which one you think applies here.

OpenStudy (anonymous):

the first one ??

myininaya (myininaya):

Simplify means they don't want any negative exponents... Which law will help us do this And no we can't use the first one because the bases aren't the same.

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

??

myininaya (myininaya):

Do any of the laws above have negative exponents?

OpenStudy (anonymous):

the third and fourth

myininaya (myininaya):

Can we use the fourth to rewrite that part that has a negative exponent?

OpenStudy (anonymous):

. yea

OpenStudy (anonymous):

can u help meeeeeeeeeee?? u still here

OpenStudy (anonymous):

@sirm3d @Daish96 @Daish96 @Dido525 @Directrix @whpalmer4 @walters @wtoliver can u help me

OpenStudy (anonymous):

@Anders96 @anabella97 @april115 @Anders96 @experimentX @experimentX @evansda can u help me

OpenStudy (anonymous):

Hollaa

OpenStudy (anonymous):

You can't add the exponents unless the a's and b's are stated to be constants.

OpenStudy (anonymous):

...self

Parth (parthkohli):

Remember that for example\[\large x^5 y^{-5} = \dfrac{x^5}{y^5}\]

Parth (parthkohli):

Can you simplify it similarly?